If anyone can help me...
The width of a rectangle is 4ft. less than the length. The area is 5ft^2. Find the length and the width.
How do I do this?

Found 2 solutions by checkley77, nerdybill:
Answer by checkley77(12844)

W=L-4
AREA=L*W
5=L(L-4)
5=L^2-4L
L^2-4L-5=0
(L-5)(L+1)=0
L-5=0
L=5 FOR THE LENGTH
5-4=1 FOR THE WIDTH.
pROOF:
5=5*1
5=5

Answer by nerdybill(7384)

The width of a rectangle is 4ft. less than the length. The area is 5ft^2. Find the length and the width.
.
Let x = length of rectangle
then because "width of a rectangle is 4ft. less than the length" we get:
x-4 = width of rectangle
.
Now, we also should know that the area of any rectangle is:
length * width
Therefore, we get out equation:
x(x-4) = 5
x^2 -4x = 5
x^2 -4x - 5 = 0
.
Factoring we get:
(x-5)(x+1) = 0
x = {-1, 5}
Since we can't have a negative length the answer must be:
x = 5 feet (length)
.
Previously, we said:
x-4 = width of rectangle
x-4 = 5-4 = 1 feet (width)